# shapes, arrows.meta, positioning Chorus: Foundation Models for Unified Data Discovery and Exploration

### v1

@@ 6. Related Work
The seminal early work in this area is WebTables (Cafarella et al., 2008). WebTables aimed to extract relational tables from messy web data into a central repository, annotating them with predicted metadata to increase discoverability. This work introduced a constellation of related tasks: schema auto-completion, attribute synonym finding, and join-graph traversal. Additional work on wrapper induction (Kushmerick et al., 1997) focused on developing shims to extracting table from heterogeneous sources.
@@
The currently deployed generation of approaches has focused on representation learning. These include turl (Deng et al., 2020) and tabert (Yin et al., 2020). Both explore the use of fine-tuned language models for similar tasks. Other systems include Doduo (Suhara et al., 2022) and TABBIE (Iida et al., 2021). Prior to these table-embedding approaches, the prior generation of data tools involved deep learning for specific tasks. The standard-bearer for this approach is Sherlock (Hulsebos et al., 2019), a tool specialized for column-type detection which utilizes a deep neural network with about input features and trained over millions of examples. Sato (Zhang et al., 2020) is another example of this approach.
Recent tutorials (Nargesian et al., 2019) outline the prevalence of the problem of unstructured document data management. A user-study of scientists conclude that “current systems fail to sufficiently support scientists in their data-seeking process” (Papenmeier et al., 2021). A dataset search survey (Chapman et al., 2020) in vldb, highlights the main open problems in this field: more natural query languages, better data integration, incorporating external knowledge, and interactive result presentation. Foundation models hold the promise of helping address many of these tasks.
Industry interest in this field is also keen. A large number of commercial solutions for data warehousing and data lakes are available. Commercial products derived from the original WebTables vision are described in the authors’ follow-up paper (Cafarella et al., 2018). Products including Amazon RedShift, Microsoft Azure Data Lake, Databricks Lakehouse are some of the commercial products in this space. Products with a more narrow focus, such as Trifacta Wrangler (Trifacta, 2023), Tableau Prep and Paxata, incorporate substantial data discovery components. Industry-led prototype systems for data discovery include Sigma Computing’s WarpGate (Cong et al., 2022) and Google Research’s Dataset Search (Brickley et al., 2019). More recently, startups such as Numbers Station raised substantial funding for fusing foundation models with enterprise data analytics. Gartner estimates the data warehouse market size at $22 billion usd.
@@ 7. Conclusion
In this work, we investigated foundation models for data discovery and exploration. We propose chorus, a system for integrating foundation models into data discovery tasks and show it provides superior performance on three example tasks: table-class annotation, column-type detection and join-column prediction. We find that chorus is more robust than prior representation-learning approaches on a variety of datasets and that its performance advantage cannot be attributed to dataset contamination. We conclude that foundation models provide a promising future as a core component of the next generation of data discovery and exploration systems.
@@ Acknowledgements.
Our thanks to Yejin Choi for guidance on and productive discussions about language models. We are grateful to the authors of TaBERT, Sherlock and DoDuo for high-quality, replicable experiment code and documentation. We appreciate Cynthia Richey’s and Kyle Deeds’ assistance with the manuscript. Finally, thanks to Magdalena Balazinska for counsel during manuscript writing.References
@@
- Corp (2018) SHI International Corp. Nvidia tesla v100 - gpu computing processor, 2018. URL https://web.archive.org/web/20230304032322/https://www.shi.com/product/34625444/NVIDIA-Tesla-V100-GPU-computing-processor.
@@
- Kandpal et al. (2022) Nikhil Kandpal, Eric Wallace, and Colin Raffel. Deduplicating training data mitigates privacy risks in language models. In Kamalika Chaudhuri, Stefanie Jegelka, Le Song, Csaba Szepesvári, Gang Niu, and Sivan Sabato, editors, International Conference on Machine Learning, ICML 2022, 17-23 July 2022, Baltimore, Maryland, USA, volume 162 of Proceedings of Machine Learning Research, pages 10697–10707. PMLR, 2022. URL https://proceedings.mlr.press/v162/kandpal22a.html.
@@
- Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, Aurélien Rodriguez, Armand Joulin, Edouard Grave, and Guillaume Lample. Llama: Open and efficient foundation language models. CoRR, abs/2302.13971, 2023. doi: 10.48550/arXiv.2302.13971. URL https://doi.org/10.48550/arXiv.2302.13971.

### v3

@@ 6. Related Work
The seminal early work is WebTables (Cafarella et al., 2008), which extracts relational tables from web data, annotated with metadata for discoverability. This work introduced the related tasks: schema auto-completion, attribute synonym finding, and join-graph traversal. Early work on wrapper induction (Kushmerick et al., 1997) also extracted tables from heterogeneous sources.
@@
The currently deployed generation of approaches has focused on representation learning. These include turl (Deng et al., 2020), tabert (Yin et al., 2020), DoDuo (Suhara et al., 2022) and tabbie (Iida et al., 2021). These explore the use of fine-tuned language models for similar tasks. Prior to these table-embedding approaches, the prior generation of data tools involved data-intensive deep learning for specific tasks, e.g. Sherlock (Hulsebos et al., 2019) and Sato (Zhang et al., 2020).
Data discovery within data lakes is an active area of research, with recent works including: unionability search (Khatiwada et al., 2023), joinability search (Zhu et al., 2019), new index structures for faster correlated dataset search (Santos et al., 2022) and end-to-end systems for data ingestion and profiling (Castelo et al., 2021). Recent tutorials (Nargesian et al., 2019; Fan et al., 2023) outline the prevalence of the problem of unstructured document data management. A user-study of scientists conclude that “current systems fail to sufficiently support scientists in their data-seeking process” (Papenmeier et al., 2021). One dataset-search survey (Chapman et al., 2020) highlights key open problems: more natural query languages, better data integration, and incorporating external knowledge.
Commercial products derived from this area of research include and include Microsoft Azure Data Lake, Databricks Lakehouse (Cafarella et al., 2018), Trifacta Wrangler (Trifacta, 2023), Tableau Prep and Paxata. Industry-led prototypes include Sigma Computing’s WarpGate (Cong et al., 2022) and Google Research’s Dataset Search (Brickley et al., 2019).
@@ 7. Conclusion
We propose chorus to integrate foundation models for data discovery. We show it provides superior performance on three exemplars: table-class annotation, column-type detection and join-column prediction. Experimentally, we find Chorus more robust than prior representation-learning approaches on a variety of benchmark datasets. We conclude that foundation models hold promise as a core component of next generation data discovery systems.
@@ Acknowledgements.
Our thanks to Yejin Choi for guidance on and productive discussions about language models. We are grateful to the authors of TaBERT, Sherlock and DoDuo for high-quality, replicable experiment code and documentation. We appreciate Cynthia Richey’s and Kyle Deeds’ assistance with the manuscript. Finally, thanks to Magdalena Balazinska for counsel during manuscript writing. This material is based upon work supported by the National Science Foundation under Grants NSF-BSF 2109922 and NSF IIS 2314527.References
@@
- Castelo et al. (2021) Sonia Castelo, Rémi Rampin, Aécio S. R. Santos, Aline Bessa, Fernando Chirigati, and Juliana Freire. Auctus: A dataset search engine for data discovery and augmentation. Proc. VLDB Endow., 14(12):2791–2794, 2021. doi: 10.14778/3476311.3476346. URL http://www.vldb.org/pvldb/vol14/p2791-castelo.pdf.
@@
- Dziri et al. (2023) Nouha Dziri, Ximing Lu, Melanie Sclar, Xiang Lorraine Li, Liwei Jiang, Bill Yuchen Lin, Peter West, Chandra Bhagavatula, Ronan Le Bras, Jena D. Hwang, Soumya Sanyal, Sean Welleck, Xiang Ren, Allyson Ettinger, Zaïd Harchaoui, and Yejin Choi. Faith and fate: Limits of transformers on compositionality. CoRR, abs/2305.18654, 2023. doi: 10.48550/arXiv.2305.18654. URL https://doi.org/10.48550/arXiv.2305.18654.
- Fan et al. (2023) Grace Fan, Jin Wang, Yuliang Li, and Renée J. Miller. Table discovery in data lakes: State-of-the-art and future directions. In Sudipto Das, Ippokratis Pandis, K. Selçuk Candan, and Sihem Amer-Yahia, editors, Companion of the 2023 International Conference on Management of Data, SIGMOD/PODS 2023, Seattle, WA, USA, June 18-23, 2023, pages 69–75. ACM, 2023. doi: 10.1145/3555041.3589409. URL https://doi.org/10.1145/3555041.3589409.
@@
- Green et al. (2007) Todd J. Green, Grigoris Karvounarakis, and Val Tannen. Provenance semirings. In Proceedings of the Twenty-Sixth ACM SIGMOD-SIGACT-SIGART Symposium on Principles of Database Systems, PODS ’07, pages 31–40, New York, NY, USA, 2007. Association for Computing Machinery. ISBN 9781595936851.
@@
- Hulsebos et al. (2021) Madelon Hulsebos, Çagatay Demiralp, and Paul Groth. Gittables: A large-scale corpus of relational tables. CoRR, abs/2106.07258, 2021. URL https://arxiv.org/abs/2106.07258.
@@
- Kayali et al. (2023) Moe Kayali, Anton Lykov, Ilias Fountalis, Nikolaos Vasiloglou, Dan Olteanu, and Dan Suciu. CHORUS: foundation models for unified data discovery and exploration. CoRR, abs/2306.09610, 2023. doi: 10.48550/arXiv.2306.09610. URL https://doi.org/10.48550/arXiv.2306.09610.
- Khatiwada et al. (2023) Aamod Khatiwada, Grace Fan, Roee Shraga, Zixuan Chen, Wolfgang Gatterbauer, Renée J. Miller, and Mirek Riedewald. SANTOS: relationship-based semantic table union search. Proc. ACM Manag. Data, 1(1):9:1–9:25, 2023. doi: 10.1145/3588689. URL https://doi.org/10.1145/3588689.
@@
- Langer et al. (1978) Ellen J Langer, Arthur Blank, and Benzion Chanowitz. The mindlessness of ostensibly thoughtful action: The role of” placebic” information in interpersonal interaction. Journal of personality and social psychology, 36(6):635, 1978.
@@
- Olsson et al. (2022) Catherine Olsson, Nelson Elhage, Neel Nanda, Nicholas Joseph, Nova DasSarma, Tom Henighan, Ben Mann, Amanda Askell, Yuntao Bai, Anna Chen, Tom Conerly, Dawn Drain, Deep Ganguli, Zac Hatfield-Dodds, Danny Hernandez, Scott Johnston, Andy Jones, Jackson Kernion, Liane Lovitt, Kamal Ndousse, Dario Amodei, Tom Brown, Jack Clark, Jared Kaplan, Sam McCandlish, and Chris Olah. In-context learning and induction heads. CoRR, abs/2209.11895, 2022. doi: 10.48550/ARXIV.2209.11895. URL https://doi.org/10.48550/arXiv.2209.11895.
@@
- Santos et al. (2022) Aécio S. R. Santos, Aline Bessa, Christopher Musco, and Juliana Freire. A sketch-based index for correlated dataset search. In 38th IEEE International Conference on Data Engineering, ICDE 2022, Kuala Lumpur, Malaysia, May 9-12, 2022, pages 2928–2941. IEEE, 2022. doi: 10.1109/ICDE53745.2022.00264. URL https://doi.org/10.1109/ICDE53745.2022.00264.
@@
- Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, Aurélien Rodriguez, Armand Joulin, Edouard Grave, and Guillaume Lample. Llama: Open and efficient foundation language models. CoRR, abs/2302.13971, 2023a. doi: 10.48550/arXiv.2302.13971. URL https://doi.org/10.48550/arXiv.2302.13971.
- Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, Dan Bikel, Lukas Blecher, Cristian Canton-Ferrer, Moya Chen, Guillem Cucurull, David Esiobu, Jude Fernandes, Jeremy Fu, Wenyin Fu, Brian Fuller, Cynthia Gao, Vedanuj Goswami, Naman Goyal, Anthony Hartshorn, Saghar Hosseini, Rui Hou, Hakan Inan, Marcin Kardas, Viktor Kerkez, Madian Khabsa, Isabel Kloumann, Artem Korenev, Punit Singh Koura, Marie-Anne Lachaux, Thibaut Lavril, Jenya Lee, Diana Liskovich, Yinghai Lu, Yuning Mao, Xavier Martinet, Todor Mihaylov, Pushkar Mishra, Igor Molybog, Yixin Nie, Andrew Poulton, Jeremy Reizenstein, Rashi Rungta, Kalyan Saladi, Alan Schelten, Ruan Silva, Eric Michael Smith, Ranjan Subramanian, Xiaoqing Ellen Tan, Binh Tang, Ross Taylor, Adina Williams, Jian Xiang Kuan, Puxin Xu, Zheng Yan, Iliyan Zarov, Yuchen Zhang, Angela Fan, Melanie Kambadur, Sharan Narang, Aurélien Rodriguez, Robert Stojnic, Sergey Edunov, and Thomas Scialom. Llama 2: Open foundation and fine-tuned chat models. CoRR, abs/2307.09288, 2023b. doi: 10.48550/arXiv.2307.09288. URL https://doi.org/10.48550/arXiv.2307.09288.
@@
- Zheng et al. (2023) Lianmin Zheng, Wei-Lin Chiang, Ying Sheng, Siyuan Zhuang, Zhanghao Wu, Yonghao Zhuang, Zi Lin, Zhuohan Li, Dacheng Li, Eric P. Xing, Hao Zhang, Joseph E. Gonzalez, and Ion Stoica. Judging llm-as-a-judge with mt-bench and chatbot arena. CoRR, abs/2306.05685, 2023. doi: 10.48550/arXiv.2306.05685. URL https://doi.org/10.48550/arXiv.2306.05685.
- Zhu et al. (2019) Erkang Zhu, Dong Deng, Fatemeh Nargesian, and Renée J. Miller. JOSIE: overlap set similarity search for finding joinable tables in data lakes. In Peter A. Boncz, Stefan Manegold, Anastasia Ailamaki, Amol Deshpande, and Tim Kraska, editors, Proceedings of the 2019 International Conference on Management of Data, SIGMOD Conference 2019, Amsterdam, The Netherlands, June 30 - July 5, 2019, pages 847–864. ACM, 2019. doi: 10.1145/3299869.3300065. URL https://doi.org/10.1145/3299869.3300065.